# MAAT: Mamba Adaptive Anomaly Transformer with association discrepancy for time series

### v1

@@ Abstract
Anomaly detection in time series poses a critical challenge in industrial monitoring, environmental sensing, and infrastructure reliability, where accurately distinguishing anomalies from complex temporal patterns remains an open problem. While existing methods, such as the Anomaly Transformer leveraging multi-layer association discrepancy between prior and series distributions and DCdetector employing dual-attention contrastive learning, have advanced the field, critical limitations persist. These include sensitivity to short-term context windows, computational inefficiency, and degraded performance under noisy and non-stationary real-world conditions. To address these challenges, we present MAAT (Mamba Adaptive Anomaly Transformer), an enhanced architecture that refines association discrepancy modeling and reconstruction quality for more robust anomaly detection. Our work introduces two key contributions to the existing Anomaly transformer architecture: Sparse Attention, which computes association discrepancy more efficiently by selectively focusing on the most relevant time steps. This reduces computational redundancy while effectively capturing long-range dependencies critical for discerning subtle anomalies. Additionally, a Mamba-Selective State Space Model (Mamba-SSM) is integrated into the reconstruction module. A skip connection bridges the original reconstruction and the Mamba-SSM output, while a Gated Attention mechanism adaptively fuses features from both pathways. This design dynamically balances fidelity and contextual enhancement, improving anomaly localization and overall detection performance.
Extensive experiments on benchmark datasets demonstrate that MAAT significantly outperforms prior methods, achieving superior anomaly distinguishability and generalization across diverse time series applications. By addressing the limitations of existing approaches, MAAT sets a new standard for unsupervised time series anomaly detection in real-world scenarios.
@@
The Mamba Adaptive Anomaly Transformer (MAAT) introduces a Sparse Attention mechanism and integrates the Mamba-Selective State Space Model (Mamba-SSM), enabling it to effectively capture both short-term and long-term temporal dependencies in time series data
MAAT outperforms existing models like the Anomaly Transformer and DCdetector by achieving higher Accuracy, precision, recall, and F1-scores across benchmark datasets.
@@ 1 Introduction
Anomaly detection in time series data is crucial across various domains, including finance, healthcare, industrial monitoring, and cybersecurity. Identifying unexpected or abnormal events in sequential data is essential for preventing system failures, detecting fraudulent activities, and ensuring operational efficiency. Traditionally, this task has been addressed using statistical methods, such as ARIMA [1] and Gaussian Processes [2], which rely on the assumption that anomalies can be identified as deviations from predicted values. However, these methods often struggle with real-world time series data complex, high-dimensional nature, especially when the underlying patterns are non-linear or when the anomalies are subtle [3].
With the advent of machine learning, more sophisticated models have been developed to address these challenges. Techniques such as Support Vector Machines (SVMs)[4], Random Forests[5], and Hidden Markov Models (HMMs)[6] were among the first machine learning methods applied to anomaly detection in time series. These models improved traditional statistical methods by learning from data rather than relying solely on predefined rules. However, they still required significant manual feature engineering and often did not fully capture the temporal dependencies inherent in sequential data.
The introduction of deep learning models marked a significant shift in time series anomaly detection. Recurrent Neural Networks (RNNs)[7], and Long Short-Term Memory (LSTM)[8] networks, in particular, have become popular due to their ability to model long-term dependencies in time series data. These models can learn complex temporal patterns directly from the data, making them well-suited for tasks such as anomaly detection, where subtle deviations from the norm need to be identified [6]. Additionally, the use of autoencoders and their variants, such as Variational Autoencoders (VAEs)[9] and Adversarial Autoencoders[10], has further advanced the field by enabling unsupervised anomaly detection through the reconstruction of standard data patterns [11].
Reconstruction-based methods are widely adopted in time series anomaly detection because they can model standard data patterns and identify deviations as anomalies. These methods typically involve training neural networks, such as autoencoders[8], to reconstruct input data, assuming that standard data can be accurately reconstructed while anomalies yield higher reconstruction errors. Autoencoders have been extensively used for this purpose because of their capacity to learn compressed representations and effectively reconstruct signals [12]. Similarly, recurrent neural networks (RNNs) and long short-term memory (LSTM) networks have been employed to capture temporal dependencies in sequential data during reconstruction tasks [8]. However, one limitation of these methods is their reliance on reconstruction loss as the sole anomaly detection criterion. This may lead to false positives when specific anomalous points exhibit patterns similar to standard data. To address this issue, The application of Transformer architectures to time series data has opened new possibilities. Transformers[13], initially developed for natural language processing, use a self-attention mechanism that allows them to model complex dependencies over long sequences. This capability is particularly advantageous for time series anomaly detection, where capturing local and global temporal patterns is crucial. In addition to the standard self-attention mechanism, various adaptations have been proposed to enhance the modeling of time series data. Gated attention mechanisms[14], for instance, integrate gating structures to dynamically control the flow of information, enabling the model to focus on relevant temporal features and improve performance in time series forecasting tasks. Sparse attention[15] mechanisms have also been introduced to improve the efficiency of Transformers when handling long sequences. By limiting the attention computation to a subset of the input sequence, sparse attention reduces computational complexity while maintaining the ability to capture essential dependencies. This approach is particularly beneficial for long-sequence time-series forecasting, where traditional attention mechanisms may become computationally prohibitive.
Recent advancements in time series anomaly detection have led to sophisticated approaches like the Anomaly Transformer [16], which leverage association discrepancy learning to distinguish normal from abnormal points without solely relying on reconstruction loss. This innovative framework employs a Gaussian kernel and attention mechanisms to model both prior-association and series-association discrepancies, comparing the expected and observed temporal associations to set a new standard in Accuracy and robustness. It addresses long-standing computational efficiency and scalability challenges that have traditionally affected reconstruction-based methods, particularly in high-dimensional or multivariate datasets. This evolution from simple statistical models to advanced deep learning architectures underscores the growing complexity and critical importance of detecting anomalies that may signal crucial events or failures. It highlights the need for models capable of pushing the boundaries of anomaly detection in increasingly complex data environments. This approach achieves state-of-the-art performance by amplifying the distinguishability between normal and abnormal data points through a minimax strategy. However, despite its advancements, the Anomaly Transformer faces several limitations. Reliance on self-attention mechanisms restricts its ability to model long-range dependencies in small window sizes, and noise and non-stationary patterns can distort association discrepancy metrics, increasing false positives during transient but normal system fluctuations. Although more efficient than traditional reconstruction-based methods.
Furthermore, DCdetector [17] simplifies the anomaly detection process by employing a dual-attention contrastive structure that eliminates the need for complex components like Gaussian kernels or reconstruction losses. Instead, it relies on contrastive learning to enhance the separation between anomalies and standard points. This approach introduces a purely contrastive learning framework for time series anomaly detection, avoiding reconstruction-based methods that often struggle with anomalies and encoder-decoder design challenges. DCdetector utilizes a dual-channel architecture, comprising patch-wise and in-patch branches, to capture both global and local temporal dependencies, ensuring robust representations for expected points while highlighting discrepancies in anomalous ones. Its channel independence patching mechanism efficiently processes multivariate time series inputs, reducing parameter complexity and mitigating overfitting risks. The model training relies on a contrastive loss function based on Kullback-Leibler divergence, which maximizes representation consistency for expected points across branches while amplifying inconsistencies for anomalies. An asymmetric branch design further prevents model collapse by maintaining distinct center vectors across branches, enabling robust self-supervised learning. This framework enhances generalization to unseen anomalies by focusing on latent pattern differences rather than prior knowledge of anomaly characteristics. However, DCdetectors suffer from several problems. One is the sensitivity to contrastive sample quality, which degrades performance on imbalanced or poorly separable data. In addition to that, the high computational overhead from pairwise comparisons and dual-branch parallelism remain issues.
Building upon these foundations, we propose MAAT (Mamba Adaptive Anomaly Transformer with association discrepancy for time series) for anomaly detection in time series. Our approach introduces a novel block architecture that incorporates skip connections and gating mechanisms in association with the mamba block in the skip connection to improve the reconstruction capability of the anomaly transformer architecture. By combining these architectural enhancements with sparse attention mechanisms, Anomaly Transformer association modeling principles, MAAT achieves superior performance in detecting anomalies across diverse datasets.
The Key contributions of this work are summarized as follows:
@@
The subsequent sections of this document are structured as follows: Section 2 presents a comprehensive overview of relevant works, including literature and methodologies, related to time series anomaly detection. Section 3 delves into our proposed MAAT architecture, detailing its unique features and potential benefits. In Section 4, we conduct Experiments, perform data analysis and provide a thorough evaluation of our model. Lastly, in Section 5, we draw insightful conclusions based on the following: In our experimentation, we compare our approach with prior methodologies and articulate the implications of our findings. This structure ensures a coherent and comprehensive understanding of our innovative methodology for unsupervised anomaly detection in time series data.
@@ 2 Related works
Time series anomaly detection has significantly advanced over the past few decades, progressing from traditional statistical models to cutting-edge deep learning and Transformer-based architectures. This section provides a detailed review of the evolution of these methods, with a particular focus on recent developments. Early time series anomaly detection work was dominated by classical statistical models such as ARIMA and Gaussian Processes (GP). ARIMA was widely used for detecting anomalies as deviations from expected values based on assumptions of linearity and stationarity [3]. However, these models were ill-suited to handle non-linear and complex patterns in real-world data. Gaussian Processes added a probabilistic framework, allowing uncertainty quantification in predictions [rasmussen2003gaussian]. Despite these improvements, ARIMA and GP struggled with high-dimensional and multivariate data, limiting their effectiveness in more complex financial and industrial monitoring applications. The rise of machine learning models has introduced more flexible approaches. Support Vector Machines (SVMs), particularly the One-Class SVM (OC-SVM), became popular for novelty detection, learning a decision boundary that encloses expected data points while flagging outliers as anomalies [4]. Despite improving over traditional statistical methods, SVMs required extensive feature engineering and could not capture temporal dependencies inherent in sequential data. Random Forests and Isolation Forests introduced new paradigms by leveraging decision trees to detect anomalies. Random Forests required fewer assumptions and handled high-dimensional data better than SVMs, while Isolation Forests identified anomalies by isolating points through recursive partitioning [18]. Both methods, however, focused on pointwise anomalies and were not optimized for capturing time dependencies in sequential data. Probabilistic methods like Hidden Markov Models (HMMs) introduced temporal dynamics into anomaly detection. HMMs model a system transitioning between hidden states over time, detecting anomalies as deviations in state transitions [5]. While HMMs represented a significant step forward, they were still limited in handling non-linear patterns and required substantial computational resources for larger datasets. The emergence of deep learning marked a significant shift in time series anomaly detection. Recurrent Neural Networks (RNNs), and more specifically, Long Short-Term Memory (LSTM) networks, became popular due to their ability to capture long-term dependencies in time series data [6]. These models excelled in detecting anomalies in healthcare (e.g., ECG data) and finance [19], where subtle deviations from standard patterns need to be identified. However, RNNs and LSTMs face limitations in scalability and can struggle with vanishing gradient problems when processing long sequences. Autoencoders (AEs), particularly Variational Autoencoders (VAEs), became central to unsupervised anomaly detection by reconstructing standard data patterns and flagging high-reconstruction errors as anomalies [11]. Adversarial Autoencoders (AAEs) and Generative Adversarial Networks (GANs) introduced adversarial training, improving the robustness of anomaly detection models. GANs, in particular, used a generator and discriminator framework to detect anomalies based on how well the model could generate realistic data samples. However, GANs and AEs both require large amounts of data for training and can suffer from instability, making them challenging to apply in certain domains [20, 21]. More recently, Transformer-based models have demonstrated significant potential in time series anomaly detection, driven by their ability to capture long-range dependencies through self-attention mechanisms. Initially developed for natural language processing tasks, transformers have been successfully applied to time series analysis without the need for recurrent structures like those in RNNs or LSTMs. The self-attention mechanism allows Transformers to capture local and global dependencies within time series data, enabling them to detect various anomaly types [13]. One prominent model is the Anomaly Transformer, which introduced the novel concept of Association Discrepancy to identify anomalies by comparing expected associations in time series data with observed ones. Using a minimax strategy, the Anomaly Transformer improves the distinguishability between normal and abnormal points, outperforming previous models across various datasets [16]. Its success highlights the growing importance of attention mechanisms in detecting complex temporal patterns, especially in high-dimensional datasets such as those found in industrial or sensor data. Several Transformer variants have emerged in recent years, each building on the strengths of the original Transformer architecture. AnomalyBERT is a model that uses self-supervised learning and a data degradation scheme to synthesize outliers and train the model to detect them. By simulating various types of anomalies, AnomalyBERT improves its ability to generalize to unseen anomalies without relying on labeled datasets [22]. Similarly, the Denoising Diffusion Mask Transformer (DDMT) combines denoising diffusion models with Transformers to enhance anomaly detection in multivariate time series, particularly in noisy environments. DDMT employs a masking mechanism to prevent information leakage, significantly improving performance in noisy industrial datasets [23]. The ongoing evolution of these models has also addressed the challenge of multivariate anomaly detection, where multiple time series are analyzed simultaneously to detect correlated anomalies. Models like Informer [24] and multi-task Transformers have been applied to multivariate settings, leveraging attention mechanisms to model dependencies between multiple variables effectively [25]. Hybrid models that combine the strengths of statistical models and deep learning are also emerging, offering improved interpretability while retaining the flexibility of neural networks. Early approaches to unsupervised time series anomaly detection primarily relied on classical statistical methods such as ARIMA models and hypothesis tests, laying the groundwork for distinguishing normal from abnormal behavior. However, these traditional methods were often limited by their inability to capture the complex interdependencies and non-stationarities inherent in high-dimensional, multivariate time series data. As the field progressed, deep learning frameworks began to emerge. In 2021, the Anomaly Transformer [16] introduced a novel anomaly-attention mechanism that quantifies the association discrepancy between normal and abnormal points. While this approach effectively leverages self-attention to capture both local and global temporal dependencies in an unsupervised manner, its quadratic complexity poses challenges for scalability, and its underlying assumptions may not hold across all types of anomalies. Building on these advances, the DC Detector [17] employs a dual attention contrastive representation learning framework. Integrating local and global attention with a contrastive loss overcomes some of the pitfalls of reconstruction-based methods. Nonetheless, the additional architectural complexity and the need for extensive hyperparameter tuning remain nontrivial hurdles. Concurrently, selective state space models have gained traction for their efficiency. Models like MAMBA [26] use a selective scanning mechanism to model long-range dependencies linearly, making them attractive for real-time and large-scale applications. However, the trade-off is a potential flexibility reduction when modeling highly non-linear interactions. Despite these significant advances, several challenges persist. One major issue is concept drift, the phenomenon where the underlying distribution of data evolves. This drift renders pre-trained models ineffective and highlights the need for adaptive learning techniques that can update dynamically as new data becomes available. In addition, scalability remains a critical challenge, especially when dealing with high-frequency time series data or real-time anomaly detection in large-scale systems. Recent efforts in developing memory-efficient Transformers and distributed learning approaches [27, 28] have begun to address these scalability issues, marking promising steps toward more adaptive and robust anomaly detection systems. Overall, the evolution of unsupervised time series anomaly detection reflects a broader trajectory across complex domains. The field is steadily progressing toward more nuanced, robust, and adaptable solutions, from early statistical methods to modern deep learning architectures that harness attention mechanisms, contrastive learning, and efficient state space representations. These advancements enhance our ability to capture intricate temporal dependencies and subtle deviations without relying on labeled data and pave the way for deploying these models in real-world, dynamic environments.

### v2

@@ Abstract
Anomaly detection in time series poses a critical challenge in industrial monitoring, environmental sensing, and infrastructure reliability, where accurately distinguishing anomalies from complex temporal patterns remains an open problem. While existing methods, such as the Anomaly Transformer leveraging multi-layer association discrepancy between prior and series distributions and DCdetector employing dual-attention contrastive learning, have advanced the field, critical limitations persist. These include sensitivity to short-term context windows, computational inefficiency, and degraded performance under noisy and non-stationary real-world conditions. To address these challenges, we present MAAT (Mamba Adaptive Anomaly Transformer), an enhanced architecture that refines association discrepancy modeling and reconstruction quality for more robust anomaly detection. Our work introduces two key contributions to the existing Anomaly transformer architecture: Sparse Attention, which computes association discrepancy more efficiently by selectively focusing on the most relevant time steps. This reduces computational redundancy while effectively capturing long-range dependencies critical for discerning subtle anomalies. A Mamba-Selective State Space Model (Mamba-SSM) is also integrated into the reconstruction module. A skip connection bridges the original reconstruction and the Mamba-SSM output, while a Gated Attention mechanism adaptively fuses features from both pathways. This design balances fidelity and contextual enhancement dynamically, improving anomaly localization and overall detection performance. Extensive experiments on benchmark datasets demonstrate that MAAT significantly outperforms prior methods, achieving superior anomaly distinguishability and generalization across diverse time series applications. By addressing the limitations of existing approaches, MAAT sets a new standard for unsupervised time series anomaly detection in real-world scenarios. You can find the link to the GitHub account in the link provided below. https://github.com/ilyesbenaissa/MAAT
@@
The Mamba Adaptive Anomaly Transformer (MAAT) features a Sparse Attention mechanism. It incorporates the Mamba-Selective State Space Model (Mamba-SSM), enabling it to effectively capture both short-term and long-term temporal dependencies in time series data.
Utilizing a block-wise sparse attention approach to minimize computational costs while enhancing anomaly detection performance.
Applying a gated attention to merge sparse attention with the outputs of the Mamba-SSM, thereby improving both anomaly localization and detection.
Leverages the State Space Model (SSM) through selective use of Mamba, facilitating more effective management of complex, noisy data.
MAAT outperforms existing models such as the Anomaly Transformer and DCdetector, achieving significantly higher F1 scores, accuracy, and recall across various benchmark datasets.
@@ 1 Introduction
Anomaly detection in time series data is critical in finance, healthcare, industrial monitoring, and cybersecurity to prevent failures, detect fraud, and ensure efficiency. Traditionally, methods like ARIMA [1] and Gaussian Processes [2] have been used to identify abnormal events in sequential data. Which rely on the assumption that anomalies can be identified as deviations from predicted values. However, these methods often struggle with real-world time series data of a complex, high-dimensional nature, especially when the underlying patterns are non-linear or when the anomalies are subtle[3]. Machine learning introduced advanced models like Support Vector Machines (SVMs)[4], Random Forests[5], and Hidden Markov Models (HMMs)[6] were among the first machine learning methods applied for time series anomaly detection, improving on traditional methods by learning from data. However, they still required extensive feature engineering and often missed complex temporal dependencies. Deep learning models such as Recurrent Neural Networks (RNNs)[7], and Long Short-Term Memory (LSTM)[8] revolutionized time series anomaly detection by capturing long-term dependencies and complex temporal patterns, ideal for identifying subtle deviations [6]. Additionally, the use of autoencoders and their variants, such as Variational Autoencoders (VAEs)[9] and Adversarial Autoencoders[10], has further advanced the field by enabling unsupervised anomaly detection through the reconstruction of standard data patterns [11]. Reconstruction-based methods, like autoencoders [8], model standard data patterns and detect anomalies through reconstruction errors, leveraging their ability to learn compressed representations [12]. Similarly, recurrent neural networks (RNNs) and long short-term memory (LSTM) networks have been employed to capture temporal dependencies in sequential data during reconstruction tasks [8]. However, reliance on reconstruction loss can cause false positives for anomalies resembling standard data. Transformers[13], with self-attention mechanisms, address this by modeling complex dependencies in time series. Enhancements like gated attention [14], for instance, integrate gating structures to dynamically control the flow of information dynamically, enabling the model to focus on relevant temporal features and improve performance in time series forecasting tasks. Sparse attention[15] mechanisms have also been introduced to improve the efficiency of Transformers when handling long sequences. By limiting the attention computation to a subset of the input sequence, sparse Attention reduces computational complexity while maintaining the ability to capture essential dependencies. This approach is particularly beneficial for long-sequence time-series forecasting, where traditional attention mechanisms may become computationally prohibitive. Recent advancements include the Anomaly Transformer [16], which uses association discrepancy learning to differentiate normal from abnormal points without relying solely on reconstruction loss. It employs Gaussian kernels and attention mechanisms to model prior- and series-association discrepancies, improving accuracy Accuracystness while addressing scalability in high-dimensional datasets. Despite achieving state-of-the-art performance through a minimax strategy, it faces limitations: self-attention struggles with long-range dependencies in small windows, and noise or non-stationary patterns can increase false positives during normal fluctuations. While more efficient than traditional methods, these challenges remain. Furthermore, DCdetector [17] simplifies anomaly detection with a dual-attention contrastive structure, eliminating complex components like Gaussian kernels or reconstruction losses. It uses contrastive learning to separate anomalies from expected points, capturing global and local dependencies through a dual-channel architecture. Its channel-independent patching reduces parameter complexity and overfitting risks, while a contrastive loss function based on Kullback-Leibler divergence enhances representation consistency. However, DCdetector faces challenges, including sensitivity to contrastive sample quality, high computational overhead from pairwise comparisons, and dual-branch parallelism. Building upon these foundations, we propose MAAT (Mamba Adaptive Anomaly Transformer with association discrepancy for time series) for anomaly detection in time series. Our approach introduces a novel block architecture that incorporates skip connections and gating mechanisms in association with the mamba block in the skip connection to improve the reconstruction capability of the anomaly transformer architecture. By combining these architectural enhancements with sparse attention mechanisms and anomaly Transformer association modeling principles, MAAT performs better in detecting anomalies across diverse datasets. The Key contributions of this work are summarized as follows:
@@
The subsequent sections of this document are structured as follows: Section 2 presents a comprehensive overview of relevant works, including literature and methodologies, related to time series anomaly detection. Section 3 delves into our proposed MAAT architecture, detailing its unique features and potential benefits. In Section 4, we conduct Experiments, perform data analysis, and provide a thorough evaluation of our model. Lastly, in Section 5, we draw insightful conclusions based on the following: In our experimentation, we compare our approach with prior methodologies and articulate the implications of our findings. This structure ensures a coherent and comprehensive understanding of our innovative methodology for unsupervised anomaly detection in time series data.
@@ 2 Related works
Time series anomaly detection has evolved from traditional statistical models like ARIMA and Gaussian Processes[3, 2]. Conventional methods like ARIMA and GP struggled with non-linear and high-dimensional data. Machine learning introduced more flexible approaches, such as (SVMs), particularly One-Class SVM (OC-SVM), which excel in novelty detection by defining a boundary around standard data and identifying outliers as anomalies [4]. Despite this improvement, SVMs could not capture temporal dependencies inherent in sequential data. Random Forests and Isolation Forests handled high-dimensional data better than SVMs, while Isolation Forests identified anomalies by isolating points through recursive partitioning [18]. However, they focused on anomalies and were not optimized for capturing time dependencies in sequential data. RNNs and LSTMs became popular due to their ability to capture long-term dependencies in time series data. These models excelled in detecting anomalies in healthcare (e.g., ECG data) and finance [19]. However, RNNs and LSTMs face limitations in scalability and can struggle with vanishing gradient problems when processing long sequences. Variational Autoencoders (VAEs) became central to unsupervised anomaly detection by reconstructing standard data patterns and flagging high-reconstruction errors as anomalies [11]. Adversarial Autoencoders (AAEs) and Generative Adversarial Networks (GANs) enhanced robustness through adversarial training. GANs use a generator-discriminator framework to detect anomalies based on realistic data generation. However, both GANs and AEs require large datasets and can face instability, limiting their application in some domains [20, 21]. Transformer-based models have recently demonstrated significant potential in time series anomaly detection by leveraging self-attention mechanisms to capture long-range dependencies [13]. Initially developed for NLP, these models eliminate the need for recurrent structures like RNNs or LSTMs, effectively modeling local and global patterns in time series data to detect anomalies. Transformer-based models have revolutionized time series anomaly detection, with innovations like the Anomaly Transformer introducing Association Discrepancy to compare expected and observed associations in data. Using a minimax strategy enhances the distinguishability of anomalies, outperforming prior methods across datasets [16]. Similarly, AnomalyBERT employs self-supervised learning and data degradation to simulate anomalies, improving generalization without labeled data [22]. The Denoising Diffusion Mask Transformer (DDMT) integrates denoising diffusion with masking mechanisms, excelling in noisy multivariate settings [23]. Multivariate anomaly detection has also advanced with models like Informer [24] and multi-task Transformers, which leverage attention mechanisms to model interdependencies among variables [25]. Hybrid models combining statistical methods with deep learning enhance interpretability while maintaining flexibility. These developments address the limitations of classical met techniques as ARIMA, which struggled with high-dimensional, non-stationary data. As the field progressed, deep learning frameworks began to emerge. In 2021, the Anomaly Transformer [16] introduced a novel anomaly-attention mechanism that quantifies the association discrepancy between normal and abnormal points. While this approach effectively leverages self-attention to capture both local and global temporal dependencies in an unsupervised manner, its quadratic complexity poses challenges for scalability, and its underlying assumptions may not hold across all types of anomalies. Building on these advances, the DC Detector [17] employs a dual attention contrastive representation learning framework. Integrating local and global Attention with a contrastive loss overcomes some of the pitfalls of reconstruction-based methods. Nonetheless, the additional architectural complexity and the need for extensive hyperparameter tuning remain nontrivial hurdles. Concurrently, selective state space models have gained traction for their efficiency. Models like MAMBA [26] use a selective scanning mechanism to model long-range dependencies linearly, making them attractive for real-time and large-scale applications. However, the trade-off is a potential flexibility reduction when modeling highly non-linear interactions. Despite these significant advances, several challenges persist. One major issue is concept drift, the phenomenon where the underlying distribution of data evolves. This drift renders pre-trained models ineffective and highlights the need for adaptive learning techniques that can update dynamically as new data becomes available. In addition, scalability remains a critical challenge, especially when dealing with high-frequency time series data or real-time anomaly detection in large-scale systems. Recent efforts in developing memory-efficient Transformers and distributed learning approaches [27, 28] have begun to address these scalability issues, marking promising steps toward more adaptive and robust anomaly detection systems. Overall, the evolution of unsupervised time series anomaly detection reflects a broader trajectory across complex domains. The field is steadily progressing toward more nuanced, robust, and adaptable solutions, from early statistical methods to modern deep learning architectures that harness attention mechanisms, contrastive learning, and efficient state space representations. These advancements enhance our ability to capture intricate temporal dependencies and subtle deviations without relying on labeled data and pave the way for deploying these models in real-world, dynamic environments.